# HalluGuard: Demystifying Data-Driven and Reasoning-Driven Hallucinations in LLMs

### v1

@@ Abstract
The reliability of Large Language Models (LLMs) in high-stakes domains such as healthcare, law, and scientific discovery is often compromised by hallucinations. These failures typically stem from two sources: data-driven hallucinations and reasoning-driven hallucinations. However, existing detection methods usually address only one source and rely on task-specific heuristics, limiting their generalization to complex scenarios. To overcome these limitations, we introduce the Hallucination Risk Bound, a unified theoretical framework that formally decomposes hallucination risk into data-driven and reasoning-driven components, linked respectively to training-time mismatches and inference-time instabilities. This provides a principled foundation for analyzing how hallucinations emerge and evolve. Building on this foundation, we introduce HalluGuard, a NTK-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard on 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, consistently achieving state-of-the-art performance in detecting diverse forms of LLM hallucinations.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed in high-stakes domains such as healthcare, law, and scientific discovery(Bommasani et al., 2021; Thirunavukarasu et al., 2023). However, adoption in these settings remains cautious, as such domains are highly regulated and demand strict compliance, interpretability, and safety guarantees(Dennstädt et al., 2025; Kattnig et al., 2024). A major barrier is the risk of hallucinations, generated content appears unfaithful or nonsensical. Such errors can have severe consequences(Dennstädt et al., 2025)—as the example in Figure 1, a generated incorrect medical diagnosis may delay treatment or lead to harmful interventions. Therefore, detecting hallucinations is not merely a technical challenge but a prerequisite for trustworthy deployment, as undetected errors undermine reliability, accountability, and user safety.
@@
Generally, hallucinations in LLMs arise from two primary sources(Ji et al., 2023; Huang et al., 2023): data-driven hallucinations, which stem from flawed, biased, or incomplete knowledge encoded during pre-training or fine-tuning; and reasoning-driven hallucinations, which originate from inference-time failures such as logical inconsistencies or breakdowns in multi-step reasoning(Zhang et al., 2023; Zhong et al., 2024). Detection methods broadly split along these two dimensions. Approaches for data-driven hallucinations often compare outputs against retrieved documents or references(Shuster et al., 2021; Min et al., 2023; Ji et al., 2023), or exploit sampling consistency as in SelfCheckGPT(Manakul et al., 2023). In contrast, methods for reasoning-driven hallucinations rely on signals of inference-time instability, including probabilistic measures such as perplexity(Ren et al., 2022), length-normalized entropy(Malinin and Gales, 2020), semantic entropy(Kuhn et al., 2023), energy-based scoring(Liu et al., 2020), and RACE(Wang et al., 2025). Others probe internal representations, for example, Inside(Chen et al., 2024a), which applies eigenvalue-based covariance metrics and feature clipping, ICR Probe(Zhang et al., 2025), which tracks residual-stream updates, and Shadows in the Attention(Wei et al., 2025), which analyzes representation drift under contextual perturbations. While these methods shed light on the mechanisms underlying hallucinations, most remain tailored to a single hallucination type and fail to capture their evolution. Yet growing evidence indicates that data-driven and reasoning-driven hallucinations often evolve during multi-step generation(Liu et al., 2025; Sun et al., 2025). As shown in Figure 1, it emerges from an initial disease misclassification and evolves into a distorted diagnosis, delaying treatments and risking fatality. This gap brings two central questions: (1) How can we develop a unified theoretical understanding of how hallucinations evolve? and (2) How can we detect them effectively and efficiently without relying on external references or task-specific heuristics?
@@
To address these challenges, we propose a unified theoretical framework–Hallucination Risk Bound, which decomposes the overall hallucination risk into two components: a data-driven term, capturing semantic deviations rooted in inaccurate, imbalanced, or noisy supervision acquired during model training; and a reasoning-driven term, reflecting instability introduced by inference-time dynamics, such as logical missteps or temporal inconsistency. This decomposition not only elucidates the mechanism behind hallucinations but also reveals how they emerge and evolve. Specifically, our analysis shows that hallucinations originate from semantic approximation gaps-captured by representational limits of the model-and are subsequently amplified by unstable rollout dynamics, evolving across decoding steps. As such, our framework offers a unified theoretical lens for characterizing the emergence and evolution of these hallucinations.
@@
Building on the theoretical foundation, we propose HalluGuard, a Neural Tangent Kernel(NTK)-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard comprehensively across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones. HalluGuard consistently achieves state-of-the-art hallucination detection performance, demonstrating its efficacy.
@@ 2 Preliminaries
There are two primary sources of hallucinations in LLMs(Ji et al., 2023; Huang et al., 2023): data-driven hallucination, which stems from incomplete or biased knowledge encoded during pre-training or fine-tuning, and reasoning-driven hallucination, which arises from unstable or inconsistent inference dynamics at decoding time. This distinction has implicitly guided a broad range of detection strategies, which we examine through these two lenses.
@@
which quantifies the average uncertainty of generating a sequence $\mathbf{y}=[y_{1},\ldots,y_{T}]$ from input $\mathbf{x}$ and $\theta$ denotes model parameters. This directly recovers Perplexity(Ren et al., 2022), where low scores imply confident predictions, while high scores indicate implausible generations due to weak priors. To capture more nuanced uncertainty, later methods extend this formulation to multi-sample settings. The Length-Normalized Entropy(Malinin and Gales, 2020) penalizes dispersion across stochastic generations $\mathcal{Y}=\{\mathbf{y}^{1},\ldots,\mathbf{y}^{K}\}$, offering a finer-grained view of model indecision. This perspective is further enriched by Semantic Entropy(Kuhn et al., 2023), which projects sampled responses into semantic space, and by energy-based scoring(Liu et al., 2020), which replaces log-probability with a learned confidence function. Collectively, these methods reflect a progression from token-level likelihoods to semantically grounded multi-sample uncertainty estimators.
@@
where $C=K\cdot(K-1)/2$, and $\text{sim}(\cdot,\cdot)$ is a similarity function such as ROUGE-L(Lin, 2004), cosine similarity, or BLEU(Chen et al., 2024b). Low scores reflect diverging generations and unstable reasoning. Several reasoning-driven detection methods can be interpreted through this lens. Early approaches used surface-level lexical overlap metrics(Lin et al., 2022b), while SelfCheckGPT(Manakul et al., 2023) advanced this by evaluating factual entailment across responses, and FActScore(Min et al., 2023) extended this further by comparing outputs to retrieved reference documents. More recent efforts probe internal signals directly: Inside(Chen et al., 2024a) analyzes the covariance spectrum of embedding representations, and RACE(Wang et al., 2025) diagnoses instability in multi-step reasoning.
@@
NTK provides a principled framework for analyzing the training dynamics in the overparameterized regime characteristic of modern LLMs(Jacot et al., 2020). Formally, for a network output $f(x,\theta)$ with input $x$ and parameters $\theta$, the NTK is defined as:
@@ 3.1 Problem Setting
Our analysis reveals that hallucination is not a unified failure mode but rather shifts with the task structure. On the instruction-following Natural benchmark(Wang et al., 2022), 88.9% of the overall 3499 errors are from logical missteps (reasoning-driven) while 11.1% are factual inaccuracies (data-driven). By contrast, on the math-focused MATH-500(Hendrycks et al., 2021), the 1985 wrong generations are dominated by 1946 reasoning errors (98.1%), with only 19 factual flaws (1.9%). This contrast highlights that, in practice, hallucinations are rarely pure but often mixtures of data-driven bias and reasoning-driven instability—motivating our formal decomposition of hallucination sources.
Problem Definition. Let $\mathcal{Y}$ denote the space of textual outputs and let $\Phi:\mathcal{Y}\rightarrow U_{h}$ be a task-specific encoder that maps textual sequences into the hypothesis space $U_{h}$, equipped with a norm $\|\cdot\|$ (e.g., task-calibrated embedding space or structured metric). We interpret each $u\in U_{h}$ as a reasoning chain, composed of step-wise logical statements. For an input $\mathbf{x}$ with ground-truth output $y^{*}\in\mathcal{Y}$, define the gold-standard reasoning chain as $u^{*}:=\Phi(y^{*})\in U_{h}$. An LLM with parameters $\theta$ emits a random sequence $Y=(Y_{1},\dots,Y_{T})$ via $p_{\theta}(y_{t}\mid y_{<t},\mathbf{x})$, yielding a predicted reasoning chain $u_{h}:=\Phi(Y)\in U_{h}$. Its expected value under the model’s decoding distribution is $\mathbb{E}[u_{h}]:=\mathbb{E}_{Y\sim p_{\theta}(\cdot\mid\mathbf{x})}[\Phi(Y)].$
We consider perturbations in a local neighborhood of the decoding process. Let $\delta\in\mathbb{R}^{r}$ parameterize a small perturbation (e.g., of the prefix tokens, step-$t$ logits, or hidden state), and let $\mathcal{B}_{\rho}:=\{\delta:\|\delta\|\leq\rho\}$. Define the perturbed decoder map $G:\mathbb{R}^{r}\!\to\!U_{h}$ by $G(\delta):=\Phi\!\big(Y(\delta)\big)$, where $Y(\delta)$ is the sequence under perturbation. Let $J\in\mathbb{R}^{d_{h}\times r}$ denote the (Gauss–Newton) Jacobian of $G$ at $\delta=0$. Our goal is to formalize how hallucination emerges and evolves in LLMs.
@@ 3.2 Hallucination Risk Bound
    $(U,\|\cdot\|)$ is a Hilbert space; $\Phi$ is measurable with unique best solution and $\|\Phi(Y)\|$ has finite second moment.
@@
    Triangle inequality holds for $\|\cdot\|$ and $\Phi$ is $L_{\Phi}$-Lipschitz w.r.t. an edit distance on $\mathcal{Y}$.
@@
    For $\delta\in\mathcal{B}_{\rho}$, the mapping $G$ admits the local expansion $G(\delta)=G(0)+J\delta+R(\delta)$, where the remainder is bounded by $\|R(\delta)\|\leq\tfrac{1}{2}H_{\star}\|\delta\|^{2}$ for some curvature constant $H_{\star}>0$.
Proposition 3.1 (Hallucination Risk Decomposition). Under A1–A3, applying the triangle inequality yields a natural split of the risk: $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\|u^{*}-\mathbb{E}[u_{h}]\|}_{\text{data-driven term}}\;+\;\underbrace{\|u_{h}-\mathbb{E}[u_{h}]\|}_{\text{reasoning-driven term}}$ This decomposition distinguishes errors caused by systematic bias in the learned representation from those introduced during stochastic rollout.
@@
Let $U_{h}\subset U$ denote the hypothesis subspace accessible to the model under perturbations. By Céa’s lemma(Céa, 1964) with curvature penalty, the data-driven term can be bounded as
@@
where $\gamma=\lambda_{\min}(\mathcal{K}_{\Phi})$ is the smallest eigenvalue of the NTK Gram matrix on embedded perturbations, and $\Lambda\leq\|T\|$ reflects the operator norm of the problem/operator mapping $\mathcal{T}$. Intuitively, the ratio $\tfrac{\Lambda}{\gamma}$ measures the conditioning of the feature map: well-conditioned NTK spectra allow a closer approximation to the true generation.
This ratio can be further controlled in terms of pretraining–finetuning mismatch:
@@
Theorem 3.2 (Hallucination Risk Bound). Let $u^{*}:=\Phi(y^{*})$ denote the semantic embedding of the ground-truth output and $u_{h}:=\Phi(Y)$ that of the model-generated output. Under Assumptions A1–A3, suppose there exists $\beta\geq 0$ such that $\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;e^{\beta T}.$ Then the total hallucination risk satisfies $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\Big(1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\cdot\tfrac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}\Big)\inf_{u\in U_{h}}\|u^{*}-u\|}_{\text{data-driven term}}\;+\;\underbrace{|\mathcal{L}|\cdot\exp\!\Big(-\tfrac{K\epsilon^{2}}{C}\Big)\cdot\alpha\big(e^{\beta T}-1\big)}_{\text{reasoning-driven term}}$

### v2

# arXiv:2601.18753v2[cs.LG] 02 Mar 2026

@@ Abstract
The reliability of Large Language Models (LLMs) in high-stakes domains such as healthcare, law, and scientific discovery is often compromised by hallucinations. These failures typically stem from two sources: data-driven hallucinations and reasoning-driven hallucinations. However, existing detection methods usually address only one source and rely on task-specific heuristics, limiting their generalization to complex scenarios. To overcome these limitations, we introduce the Hallucination Risk Bound, a unified theoretical framework that formally decomposes hallucination risk into data-driven and reasoning-driven components, linked respectively to training-time mismatches and inference-time instabilities. This provides a principled foundation for analyzing how hallucinations emerge and evolve. Building on this foundation, we introduce HalluGuard, a NTK-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard on 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, consistently achieving state-of-the-art performance in detecting diverse forms of LLM hallucinations. We open-source our proposed HalluGuard model at HalluGuard.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed in high-stakes domains such as healthcare, law, and scientific discovery (Bommasani et al., 2021; Thirunavukarasu et al., 2023; Ke et al., 2025). However, adoption in these settings remains cautious, as such domains are highly regulated and demand strict compliance, interpretability, and safety guarantees (Dennstädt et al., 2025; Kattnig et al., 2024). A major barrier is the risk of hallucinations, generated content appears unfaithful or nonsensical. Such errors can have severe consequences (Dennstädt et al., 2025), as the example in Figure 1, a generated incorrect medical diagnosis may delay treatment or lead to harmful interventions. Therefore, detecting hallucinations is not merely a technical challenge but a prerequisite for trustworthy deployment, as undetected errors undermine reliability, accountability, and user safety.
@@
Generally, hallucinations in LLMs arise from two primary sources (Ji et al., 2023; Huang et al., 2025): data-driven hallucinations, which stem from flawed, biased, or incomplete knowledge encoded during pre-training or fine-tuning; and reasoning-driven hallucinations, which originate from inference-time failures such as logical inconsistencies or breakdowns in multi-step reasoning (Zhang et al., 2023; Zhong et al., 2024). Detection methods broadly split along these two dimensions. Approaches for data-driven hallucinations often compare outputs against retrieved documents or references (Shuster et al., 2021; Min et al., 2023; Ji et al., 2023), or exploit sampling consistency as in SelfCheckGPT (Manakul et al., 2023). In contrast, methods for reasoning-driven hallucinations rely on signals of inference-time instability, including probabilistic measures such as perplexity (Ren et al., 2023), length-normalized entropy (Malinin and Gales, 2021), semantic entropy (Kuhn et al., 2023), energy-based scoring (Liu et al., 2020), and RACE (Wang et al., 2025). Others probe internal representations, for example, Inside (Chen et al., 2024), which applies eigenvalue-based covariance metrics and feature clipping, ICR Probe (Zhang et al., 2025), which tracks residual-stream updates, and Shadows in the Attention (Wei et al., 2025), which analyzes representation drift under contextual perturbations. While these methods shed light on the mechanisms underlying hallucinations, most remain tailored to a single hallucination type and fail to capture their evolution. Yet growing evidence indicates that data-driven and reasoning-driven hallucinations often evolve during multi-step generation (Liu et al., 2025; Sun et al., 2025). As shown in Figure 1, it emerges from an initial disease misclassification and evolves into a distorted diagnosis, delaying treatments and risking fatality. This gap brings two central questions: (1) How can we develop a unified theoretical understanding of how hallucinations evolve? (2) How can we detect them effectively and efficiently without relying on external references or task-specific heuristics?
@@
To address these challenges, we propose a unified theoretical framework–Hallucination Risk Bound, which decomposes the overall hallucination risk into two components: a data-driven term, capturing semantic deviations rooted in inaccurate, imbalanced, or noisy supervision acquired during model training; and a reasoning-driven term, reflecting instability introduced by inference-time dynamics, such as logical missteps or temporal inconsistency. This decomposition not only elucidates the mechanism behind hallucinations but also reveals how they emerge and evolve. Specifically, our analysis shows that hallucinations originate from semantic approximation gaps, captured by representational limits of the model, and are subsequently amplified by unstable rollout dynamics, evolving across decoding steps. As such, our framework offers a unified theoretical lens for characterizing the emergence and evolution of these hallucinations.
@@
Building on the theoretical foundation, we propose HalluGuard, a Neural Tangent Kernel(NTK)-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard comprehensively across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones. HalluGuard consistently achieves state-of-the-art hallucination detection performance, demonstrating its efficacy. We open-source our proposed HalluGuard model at HalluGuard.
@@ 2 Preliminaries
There are two primary sources of hallucinations in LLMs (Ji et al., 2023; Huang et al., 2025): data-driven hallucination, which stems from incomplete or biased knowledge encoded during pre-training or fine-tuning, and reasoning-driven hallucination, which arises from unstable or inconsistent inference dynamics at decoding time. This distinction has implicitly guided a broad range of detection strategies, which we examine through these two lenses.
@@
which quantifies the average uncertainty of generating a sequence $\mathbf{y}=[y_{1},\ldots,y_{T}]$ from input $\mathbf{x}$ and $\theta$ denotes model parameters. This directly recovers Perplexity (Ren et al., 2023), where low scores imply confident predictions, while high scores indicate implausible generations due to weak priors. To capture more nuanced uncertainty, later methods extend this formulation to multi-sample settings. The Length-Normalized Entropy (Malinin and Gales, 2021) penalizes dispersion across stochastic generations $\mathcal{Y}=\{\mathbf{y}^{1},\ldots,\mathbf{y}^{K}\}$ where $K$ denotes the number of independent stochastic rollouts sampled from the model for a given input, offering a finer-grained view of model indecision. This perspective is further enriched by Semantic Entropy (Kuhn et al., 2023), which projects sampled responses into semantic space, and by energy-based scoring (Liu et al., 2020), which replaces log-probability with a learned confidence function. Collectively, these methods reflect a progression from token-level likelihoods to semantically grounded multi-sample uncertainty estimators.
@@
where $C=K\cdot\frac{(K-1)}{2}$, and $\text{sim}(\cdot,\cdot)$ is a similarity function such as ROUGE-L (Lin, 2004), cosine similarity, or BLEU (Chen et al., 2023). Low scores reflect diverging generations and unstable reasoning. Several reasoning-driven detection methods can be interpreted through this lens. Early approaches used surface-level lexical overlap metrics (Lin et al., 2022b), while SelfCheckGPT (Manakul et al., 2023) advanced this by evaluating factual entailment across responses, and FActScore (Min et al., 2023) extended this further by comparing outputs to retrieved reference documents. More recent efforts probe internal signals directly: Inside (Chen et al., 2024) analyzes the covariance spectrum of embedding representations, and RACE (Wang et al., 2025) diagnoses instability in multi-step reasoning.
@@
NTK provides a principled framework for analyzing the training dynamics in the overparameterized regime characteristic of modern LLMs (Jacot et al., 2018). Formally, for a network output $f(x,\theta)$ with input $x$ and parameters $\theta$, the NTK is defined as:
@@ 3.1 Problem Setting
Our analysis reveals that hallucination is not a unified failure mode but rather shifts with the task structure. On the instruction-following Natural benchmark (Wang et al., 2022), 88.9% of the overall 3499 errors are from logical missteps (reasoning-driven) while 11.1% are factual inaccuracies (data-driven). By contrast, on the math-focused MATH-500 (Hendrycks et al., 2021), the 1985 wrong generations are dominated by 1946 reasoning errors (98.1%), with only 19 factual flaws (1.9%). This contrast highlights that, in practice, hallucinations are rarely pure but often mixtures of data-driven bias and reasoning-driven instability-motivating our formal decomposition of hallucination sources.
Problem Definition.
Let $\mathcal{Y}$ denote the discrete space of all possible finite-length textual token sequences. We define a continuous semantic embedding space $U_{h}\subseteq\ \mathbb{R}^{d_{h}}$ equipped with a norm $\|\cdot\|$. Each vector $u\in U_{h}$ represents the semantic representation of a reasoning chain composed of step-wise logical statements. We define a task-specific encoder $\Phi:\mathcal{Y}\rightarrow U_{h}$ that maps a discrete textual sequence into this continuous hypothesis space. In this framework, for an input $\mathbf{x}$ with a ground-truth output sequence $y^{*}\in\mathcal{Y}$, we define the target semantic representation as $u^{*}:=\Phi(y^{*})\in U_{h}.$ An LLM with parameters $\theta$ emits a random sequence $Y=(Y_{1},\dots,Y_{T})\in\mathcal{Y}$ via the autoregressive decoding distribution $p_{\theta}(y_{t}\mid y_{<t},\mathbf{x})$, yielding a predicted semantic representation $u_{h}:=\Phi(Y)\in U_{h}.$ Thus, the model’s expected sematic output is defined as $\mathbb{E}[u_{h}]:=\mathbb{E}_{Y\sim p_{\theta}(\cdot\mid\mathbf{x})}[\Phi(Y)].$
To analyze inference dynamics, we consider perturbations in a local neighborhood of the decoding process. Let $\mathbb{R}^{r}$ denote the $r$-dimensional continuous space of the model’s internal states (e.g., prefix embeddings or hidden activations). We parameterize a small perturbation by $\delta\in\mathbb{R}^{r},$ restricted to a local $\ell_{2}$-ball $\mathcal{B}_{\rho}:=\{\delta\in\mathbb{R}^{r}:\|\delta\|_{2}\leq\rho\}.$ Let $P_{\theta}(\cdot\mid\mathbf{x},\delta)$ denote the perturbed decoding distribution induced by $\delta$. We define the mean semantic response map $G_{Y}:\mathbb{R}^{r}\rightarrow U_{h},G_{Y}(\delta):=\mathbb{E}_{Y\sim P_{\theta}(\cdot\mid\mathbf{x},\delta)}[\Phi(Y)]$ with its corresponding inference Jacobian $J:=DG_{Y}(0)\in\mathbb{R}^{d_{h}\times r}$. Thus, we formally define the problem as follows:
Problem 1 (Hallucination Dynamics Characterization).
Given: (1) The target semantic representation $u^{*}:=\Phi(y^{*})\in U_{h}$ for a ground-truth output $y^{*}\in\mathcal{Y}$; (2) the random sequence $Y\in\mathcal{Y}$ emitted via the autoregressive decoding distribution $p_{\theta}(y_{t}\mid y_{<t},\mathbf{x})$, yielding a predicted representation $u_{h}:=\Phi(Y)$ with expected value $\mathbb{E}[u_{h}]$ ; and (3) the inference constraints defined by a local perturbation $\delta\in\mathbb{R}^{r}$ restricted to the $\ell_{2}$-ball $\mathcal{B}_{\rho}$
Find: A formal geometric mechanism to characterize how hallucinations emerge and evolve by analyzing the Mean Semantic Response Map $G_{Y}(\delta)$ and the Inference Jacobian $J$, which captures the sensitivity of the model’s reasoning trajectory to internal instabilities.
@@ 3.2 Hallucination Risk Bound
    $(U_{h},\|\cdot\|)$ is a finite-dimensional Hilbert space. The encoder $\Phi:\mathcal{Y}\to U_{h}$ is measurable, and the random variable $\Phi(Y)$ has finite second moment under the model’s unperturbed decoding distribution: $\mathbb{E}_{Y\sim p_{\theta}(\cdot\mid\mathbf{x})}\big[\|\Phi(Y)\|^{2}\big]<\infty.$ This ensures that the mean semantic representation $\mathbb{E}[\Phi(Y)]$ is well-defined in $U_{h}$.
@@
    Let $(\mathcal{Y},d_{\mathcal{Y}})$ be the discrete metric space equipped with edit distance. The encoder $\Phi:(\mathcal{Y},d_{\mathcal{Y}})\rightarrow(U_{h},\|\cdot\|)$ is $L_{\Phi}$-Lipschitz continuous: $\|\Phi(y)-\Phi(y^{\prime})\|\leq L_{\Phi}\,d_{\mathcal{Y}}(y,y^{\prime})\quad\forall y,y^{\prime}\in\mathcal{Y}.$
@@
    For any perturbation $\delta$ in the closed ball $\mathcal{B}_{\rho}:=\{\delta\in\mathbb{R}^{r}:\|\delta\|_{2}\leq\rho\}$, the mean semantic response map $G_{Y}(\delta)=\mathbb{E}_{Y\sim P_{\theta}(\cdot\mid\mathbf{x},\delta)}[\Phi(Y)]$ is twice Fréchet differentiable in a neighborhood of $\delta=0$ and admits the expansion $G_{Y}(\delta)=G_{Y}(0)+J\delta+R(\delta),$ where $J=DG_{Y}(0)\in\mathbb{R}^{d_{h}\times r}$ and the remainder satisfies $\|R(\delta)\|\leq\frac{1}{2}H_{\star}\|\delta\|_{2}^{2},\forall\delta\in\mathcal{B}_{\rho},$ for some constant $H_{\star}>0$.
Proposition 3.1 (Hallucination Risk Decomposition). Under A1-A3, applying the triangle inequality yields a natural split of the risk: $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\|u^{*}-\mathbb{E}[u_{h}]\|}_{\text{data-driven term}}\;+\;\underbrace{\|u_{h}-\mathbb{E}[u_{h}]\|}_{\text{reasoning-driven term}}$ This decomposition distinguishes errors caused by systematic bias in the learned representation from those introduced during stochastic rollout.
@@
Let $U_{h}$ denote the hypothesis subspace accessible to the model under perturbations. By Céa’s lemma(Céa, 1964) with curvature penalty, the data-driven term can be bounded as
@@
where $\gamma=\lambda_{\min}(\mathcal{K}_{\Phi})$ is the smallest eigenvalue of the NTK Gram matrix on embedded perturbations $\mathcal{K}_{\Phi}$, and $\Lambda\leq\|\mathcal{T}\|$, where $\mathcal{T}:U_{h}\to U_{h}$ denotes the operator mapping. Intuitively, the ratio $\tfrac{\Lambda}{\gamma}$ measures the conditioning of the feature map: well-conditioned NTK spectra allow a closer approximation to the true generation.
Thus, the ratio can be further controlled in terms of pretraining-finetuning mismatch:
@@
Theorem 3.2 (Hallucination Risk Bound). Let $u^{*}:=\Phi(y^{*})$ denote the semantic embedding of the ground-truth output and $u_{h}:=\Phi(Y)$ that of the model-generated output. Under Assumptions A1-A3, suppose there exists $\beta\geq 0$ such that $\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;e^{\beta T}.$ Then the total hallucination risk satisfies $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\Big(1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\cdot\tfrac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}\Big)\inf_{u\in U_{h}}\|u^{*}-u\|}_{\text{data-driven term}}\;+\;\underbrace{|\mathcal{L}|\cdot\exp\!\Big(-\tfrac{K\epsilon^{2}}{C}\Big)\cdot\alpha\big(e^{\beta T}-1\big)}_{\text{reasoning-driven term}}$ Here, $|\mathcal{L}|$ denotes the total sampled trajectories.